Each of the eight digits 1,2,3,4,5,6,7,8 is written on a piece of paper and all of the 8 pieces of paper are then placed in a hat. Find the probability that the number selected is
a. odd
b. greater than 6
c. odd AND greater than 5
d. odd OR greater than 5
If the same eight pieces of paper are placed in a hat, find the probability WITHOUT replacement that 2 pieces selected from the hat will
a. both be odd
b. both be less than 4

Answer by stanbon(75887)

Each of the eight digits 1,2,3,4,5,6,7,8 is written on a piece of paper and all of the 8 pieces of paper are then placed in a hat. Find the probability that the number selected is
a. odd: 4/8 = 1/2
------------------------------
b. greater than 6: 2/8 = 1/4
------------------------------
c. odd AND greater than 5: 1/8
-----------------------------------
d. odd OR greater than 5: (4/8)+(3/8)-(1/8) = 6/8 = 3/4
-----------------------------------
If the same eight pieces of paper are placed in a hat, find the probability WITHOUT replacement that 2 pieces selected from the hat will
a. both be odd: (4/8)(3/7) = 12/56 = 3/14
------
b. both be less than 4:: (3/8)(2/7) = 6/56 = 3/28
==============================
Cheers,
Stan H.